# Non-Rigid Designators in Modal and Temporal Free Description Logics

### v1

@@ Appendix A Details on Introduction and Preliminaries
${\begin{array}[]{lcl}\mathfrak{M},w\models^{\sigma}P(x_{1},\ldots,x_{n})&\text{iff}&(\sigma(x_{1}),\ldots,\sigma(x_{n}))\in P^{\mathcal{I}_{w}};\\ \mathfrak{M},w\models^{\sigma}x=y&\text{iff}&\sigma(x)=\sigma(y);\\ \mathfrak{M},w\models^{\sigma}\neg\varphi&\text{iff}&\mathfrak{M},w\not\models^{\sigma}\varphi;\\ \mathfrak{M},w\models^{\sigma}\varphi\land\psi&\text{iff}&\mathfrak{M},w\models^{\sigma}\varphi\text{ and }\mathfrak{M},w\models^{\sigma}\psi;\\ \mathfrak{M},w\models^{\sigma}\exists x\,\varphi&\text{iff}&\parbox[t][0.0pt][t]{119.50157pt}{$\mathfrak{M},w\models^{\sigma^{\prime}}\varphi$, for some $x$-variant $\sigma^{\prime}$ of $\sigma$;}\\ \mathfrak{M},w\models^{\sigma}\Diamond_{i}\varphi&\text{iff}&\parbox[t][0.0pt][t]{119.50157pt}{$\mathfrak{M},v\models^{\sigma}\varphi$, for some $v\in W$ such that $wR_{i}v$;}\\ \mathfrak{M},w\models^{\sigma}\langle\lambda x.\varphi\rangle(\tau)&\text{iff}&\parbox[t][0.0pt][t]{119.50157pt}{$\tau$ designates under $\sigma$ at $w$ of $\mathfrak{M}$ and $\mathfrak{M},w\models^{\sigma^{\prime}}\varphi$, where $\sigma^{\prime}$ is the $x$-variant of $\sigma$ with $\sigma^{\prime}(x)={\tau}^{\mathfrak{M},{w}}_{\sigma}$.}\end{array}}$
@@ Appendix B Proofs for Section 3
Clearly, for any interpretation $\mathfrak{M}^{\ast}$ such that $\mathfrak{M}^{\ast}\models\mathcal{O}^{\ast}$, we have $\mathfrak{M}^{\ast}\models\mathcal{O}$ as well.
@@ Appendix C Proofs for Section 4
$\top$ $\sqsubseteq\exists u.(\{a\}^{\sharp}\sqcap\neg\exists^{\neq}u.\{a\}^{\sharp}),\text{ for every }a\in{\sf N_{I}}\text{ in }\mathcal{O},$
@@ Appendix D Proofs for Section 5
$\vec{w}=w_{0}i_{0}w_{1}\cdots i_{m-1}w_{m},$ (7)
where $1\leq i_{j}\leq n$, $i_{j}\not=i_{j+1}$, and each $R_{i}$ is the smallest equivalence relation containing all pairs of the form $(\vec{w},\vec{w}iw)\in W\times W$. We call $w_{0}$ the root of $\mathfrak{F}$. We define the depth of $\mathfrak{F}$ as the maximal $m$ such that $W$ contains a word of the form (7). A tree-shaped $\mathbf{S5}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$-model takes the form $\mathfrak{M}=(\mathfrak{F},\Delta,\mathcal{I})$ with $\mathfrak{F}$ a tree-shaped $\mathbf{S5}^{n}$-frame.
@@
where $W^{\ast}$ is the set of words $\vec{w}$ of the form (7) with $(w_{j},w_{j+1})\in R_{i_{j}}$, $w_{j}\not=w_{j+1}$, $i_{j}\not=i_{j+1}$, and $m\leq K$, and $R^{\ast}_{i}$ is the smallest equivalence relation containing all pairs $(\vec{w},\vec{w}iw)\in W^{\ast}\times W^{\ast}$. Define a model $\mathfrak{M}^{\ast}=(\mathfrak{F}^{\ast},\Delta^{\ast},\mathcal{I}^{\ast})$ by setting $\Delta^{\ast}=\Delta$ and $\mathcal{I}^{\ast}_{\vec{w}}=\mathcal{I}_{w_{m}}$ for $\vec{w}\in W^{\ast}$ of the form (7). Clearly $\mathfrak{M}^{\ast}$ is tree-shaped. Moreover, for all $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concepts $D$ of modal depth not exceeding $K$, one can show by induction that $D^{\mathcal{I}^{\ast}_{w_{0}}}=D^{\mathcal{I}_{w_{0}}}$. In particular, $C^{\mathcal{I}^{\ast}_{w_{0}}}\neq\emptyset$. ∎
@@
    for every $w\in W$ and for every ${\boldsymbol{t}}\in\boldsymbol{q}(w),\text{ with }\{a\}\in{\boldsymbol{t}}$, there exists exactly one $\rho\in\mathfrak{R}$ such that $\{a\}\in\rho(w)$.
@@
For every $C\in\mathsf{con}(C_{0})$, $w\in W$ and $\rho\in\Delta$, $\rho\in C^{\mathcal{I}_{w}}$ iff $C\in\rho(w)$.
@@
For each $\vec{w}\in W$ and ${\boldsymbol{t}}\in\boldsymbol{q}(\vec{w})$ we fix a proto-run, $\rho_{\vec{w},{\boldsymbol{t}}}\in\mathfrak{R}$, for $\vec{w}$ and ${\boldsymbol{t}}$ with $\rho_{\vec{w},{\boldsymbol{t}}}(\vec{w})={\boldsymbol{t}}$. Let $\mathcal{F}(\vec{w})$ denote the set of selected proto-runs through $\boldsymbol{q}(\vec{w})$, that is $\mathcal{F}(\vec{w})=\{\rho_{\vec{w},{\boldsymbol{t}}}\mid{\boldsymbol{t}}\in\boldsymbol{q}(\vec{w})\}$. We also fix for any $\Diamond_{i}C\in{\boldsymbol{t}}\in\boldsymbol{q}(\vec{w})$ a proto-witness $\vec{w}^{\prime}=v_{\vec{w},i,C,{\boldsymbol{t}}}\in W$ such that $\vec{w}^{\prime}$ takes the form $\vec{w}iv$ (and so $(\vec{w},\vec{w}^{\prime})\in R_{i}$) and $C\in\rho_{\vec{w},{\boldsymbol{t}}}(\vec{w}^{\prime})$, whenever such a $\vec{w}^{\prime}$ exists. Note that such a $\vec{w}^{\prime}$ might not exist since by the definition of $R_{i}$ as equivalence closure in tree-shaped models the witness for $\Diamond_{i}C$ might also be ${\boldsymbol{t}}$ itself if $C\in{\boldsymbol{t}}$ or $\vec{w}$ could take the form $\vec{w}_{0}iw$ for some $\vec{w}_{0}$ and either $C\in\rho_{\vec{w},{\boldsymbol{t}}}(\vec{w}_{0})$ or there is $w^{\prime}$ with $\vec{w}_{0}iw^{\prime}\in W$ and $C\in\rho_{\vec{w},{\boldsymbol{t}}}(\vec{w}_{0}iw^{\prime})$.
@@
$W_{j+1}^{i}=\{v_{\vec{w},i,C,{\boldsymbol{t}}}\mid\Diamond_{i}C\in{\boldsymbol{t}}\in\boldsymbol{q}(\vec{w}),\vec{w}\in W_{j}\}.$
@@
$R_{i}^{\prime}=R_{i}\cap(W^{\prime}\times W^{\prime}),\quad\boldsymbol{q}^{\prime}(\vec{w})=\boldsymbol{q}(\vec{w})\text{ for $\vec{w}\in W^{\prime}$}$
Also let $\mathfrak{R}^{\prime}$ denote the set of restrictions of the proto-runs $\rho_{\vec{w},{\boldsymbol{t}}}$ with $\vec{w}\in W^{\prime}$ to $W^{\prime}$. Let $\mathfrak{F}^{\prime}=(W^{\prime},R_{1},^{\prime}\ldots,R_{n}^{\prime})$ and $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$. Then $\mathfrak{Q}^{\prime}$ is a quasimodel except that the proto-runs in $\mathfrak{R}^{\prime}$ might not all satisfy (R2). Note that (R2) holds for every proto-run $\rho_{\vec{w},{\boldsymbol{t}}}$ at the world $\vec{w}$ for those $\Diamond_{i}C$ for which $i_{m-1}\not=i$ (assuming that $\vec{w}$ takes the form (7)) but that there might not be witnesses for $\Diamond_{i}C$ on $\vec{w}^{\prime}\not=\vec{w}$ and if $i_{m-1}=i$. To ensure (R2) also in these cases we introduce, in a careful way, copies of existing worlds that provide the witnesses for such $\Diamond_{i}C$.
@@
Take for any $\vec{w}\in W^{\prime}$ and $\rho\in\mathfrak{R}^{\prime}$ the bijection $\sigma_{\vec{w},\rho}\in B(\mathfrak{R}^{\prime})$ that swaps $\rho$ with the proto-run $\rho_{\vec{w},{\boldsymbol{t}}}$ with ${\boldsymbol{t}}=\rho(\vec{w})$ and maps the remaining runs to themselves. The $\vec{w}$-repair set, Rep$(\vec{w})$, is the set of all $\sigma_{\vec{w},\rho}$ with $\rho\in\mathfrak{R}^{\prime}$. Note that the identity function, $id$, on $\mathfrak{R}^{\prime}$ is an element of Rep$(\vec{w})$.
For any $(\vec{w},\sigma)\in W^{\prime}\times B(\mathfrak{R}^{\prime})$, define
$Suc_{i}(\vec{w},\sigma)=\{(\vec{w}^{\prime},\sigma^{\prime})\in W^{\prime}\times B\mid\vec{w}^{\prime}=\vec{w}iw\in W^{\prime},\\ \sigma^{\prime}=\tau\circ\sigma\text{ for some }\tau\in\text{Rep}(\vec{w})\}$
@@
$\vec{u}=(\vec{w}_{0},\sigma_{0})i_{0}(\vec{w}_{1},\sigma_{1})i_{1}\cdots i_{m-1}(\vec{w}_{m},\sigma_{m}),$ (8)
where $(\vec{w}_{0},\sigma_{0}):=(w_{0},id)$ and $(\vec{w}_{i+1},\sigma_{i+1})\in Suc_{i}(\vec{w}_{i},\sigma_{i})$ for $0\leq i<m$.
Define $\boldsymbol{q}^{\prime\prime}$ by setting $\boldsymbol{q}^{\prime\prime}(\vec{u})=\boldsymbol{q}^{\prime}(\vec{w}_{m})$ for any $\vec{u}\in W^{\prime\prime}$ of the form (8). Associate with any run $\rho\in\mathfrak{R}^{\prime}$ a unique run $\rho^{\uparrow}$ on $W^{\prime\prime}$ by setting $\rho^{\uparrow}(\vec{u}):=\sigma_{m}(\rho)(\vec{w}_{m})$ for any $\vec{u}\in W^{\prime\prime}$ of the form (8) and let ′′=$\mathfrak{R}^{\prime\prime}=\{\rho^{\uparrow}\mid\rho\in\mathfrak{R}^{\prime}\}$.
@@
Observe first that $\rho\mapsto\rho^{\uparrow}$ is a bijection from $\mathfrak{R}^{\prime}$ onto $\mathfrak{R}^{\prime\prime}$. To show this observe that the mapping is surjective by definition. To show that it is injective, assume that $\rho_{1}\not=\rho_{2}$. Take $\vec{w}_{m}\in W^{\prime}$ of the form (7) with $\rho_{1}(\vec{w}_{m})\not=\rho_{2}(\vec{w}_{m})$. Then let
$\vec{v}=(\vec{w}_{0},id)i_{0}(\vec{w}_{1},id)i_{1}\cdots i_{m-1}(\vec{w}_{m},id)\in W^{\prime\prime},$
where $\vec{w}_{i+1}=\vec{w}_{i}iw_{i+1}$ for $i<m$. We have $\rho_{1}^{\uparrow}(\vec{v})=id(\rho_{1})(\vec{w}_{m})\not=id(\rho_{2})(\vec{w}_{m})$, as required.
@@
We prove (M1) and (M2) in one go by showing that for any $\vec{u}\in W^{\prime\prime}$ of the form (8) and ${\boldsymbol{t}}\in\boldsymbol{q}^{\prime\prime}(\vec{u})$ there is a bijection between $V_{0}=\{\rho\in\mathfrak{R}^{\prime\prime}\mid\rho(\vec{u})={\boldsymbol{t}}\}$ and $V_{1}=\{\rho\in\mathfrak{R}^{\prime}\mid\rho(\vec{w}_{m})={\boldsymbol{t}}\}$ (recall that $\boldsymbol{q}^{\prime\prime}(\vec{u})=\boldsymbol{q}^{\prime}(\vec{w}_{m})$).
Indeed, we claim that $F\colon V_{0}\rightarrow V_{1}$ defined by setting $F(\rho^{\uparrow})=\sigma_{m}(\rho)$ is a bijection. As $\rho\mapsto\rho^{\uparrow}$ is bijective and $\rho^{\uparrow}(\vec{u})=\sigma_{m}(\rho)(\vec{w}_{m})$, $F$ is a well defined function into $V_{1}$. It is injective since $\sigma_{m}$ is also bijective, by definition. It is surjective since for any $\rho\in\mathfrak{R}^{\prime}$ with $\rho(\vec{w}_{m})={\boldsymbol{t}}$ we have $F((\sigma_{m}^{-1}(\rho))^{\uparrow})=\sigma_{m}\sigma_{m}^{-}(\rho)=\rho$ and $\sigma_{m}^{-1}(\rho)\in\mathfrak{R}^{\prime}$.
We next show (R1) and (R2). Assume that $\Diamond_{i}C\in\mathsf{con}(C_{0})$ and $\rho^{\uparrow}\in\mathfrak{R}^{\prime\prime}$. Consider $\vec{u}\in W^{\prime\prime}$ of the form (8).
For (R1), assume there exists $\vec{v}\in W^{\prime\prime}$ such that $\vec{u}R_{i}^{\prime\prime}\vec{v}$ and $C\in\rho^{\uparrow}(\vec{v})$. We have to show that $\Diamond_{i}C\in\rho^{\uparrow}(\vec{u})$. Observe that the runs in $\mathfrak{R}^{\prime}$ satisfy (R1) in $\mathfrak{Q}^{\prime}$ and that, moreover,
@@
    if $\Diamond_{i}C\in\rho(\vec{w})$ with $\rho\in\mathfrak{R}^{\prime}$ and $(\vec{w},\vec{w}^{\prime})\in R_{i}^{\prime}$, then $\Diamond_{i}C\in\rho(\vec{w}^{\prime})$.
@@
    $\vec{v}=\vec{u}$. Then $\Diamond_{i}C\in\rho^{\uparrow}(\vec{u})$ since we have $\Diamond_{i}C\in{\boldsymbol{t}}$ whenever $C\in{\boldsymbol{t}}$ for $\Diamond_{i}C\in\mathsf{con}(C_{0})$ for any ${\boldsymbol{t}}$.
@@
    $\vec{v}=\vec{u}i(\vec{u}iw_{m+1},\sigma_{m+1})$ for some $(w_{m+1},\sigma_{m+1})$ with $\vec{w}_{m}iw_{m+1}\in W^{\prime}$ with $\sigma_{m+1}=\tau\circ\sigma_{n}$ and $\tau\in\text{Rep}(\vec{w}_{m})$.
    We have $\rho^{\uparrow}(\vec{u})=\sigma_{m}(\rho)(\vec{w}_{m})$ and $\rho^{\uparrow}(\vec{v})=\tau\circ\sigma_{m}(\rho)(\vec{w}_{m}iw_{m+1})$. From $C\in\tau\circ\sigma_{m}(\rho)(\vec{w}_{m}iw_{m+1})$ it follows that $\Diamond_{i}C\in\tau\circ\sigma_{m}(\rho)(\vec{w}_{m})$ since $\tau\circ\sigma_{m}(\rho)$ satisfies condition (R1) for runs in $\mathfrak{Q}^{\prime}$. Now $\tau\circ\sigma_{m}(\rho)$ coincides with $\sigma_{m}(\rho)$ on $\vec{w}_{m}$. Hence $\Diamond_{i}C\in\sigma_{m}(\rho)(\vec{w}_{m})$, as required.

### v2

@@ Appendix A Details on Introduction and Preliminaries
${\begin{array}[]{lcl}\mathfrak{M},w\models^{\sigma}P(x_{1},\ldots,x_{n})&\text{iff}&(\sigma(x_{1}),\ldots,\sigma(x_{n}))\in P^{\mathcal{I}_{w}};\\ \mathfrak{M},w\models^{\sigma}x_{1}=x_{2}&\text{iff}&\sigma(x_{1})=\sigma(x_{2});\\ \mathfrak{M},w\models^{\sigma}\neg\varphi&\text{iff}&\mathfrak{M},w\not\models^{\sigma}\varphi;\\ \mathfrak{M},w\models^{\sigma}\varphi\land\psi&\text{iff}&\mathfrak{M},w\models^{\sigma}\varphi\text{ and }\mathfrak{M},w\models^{\sigma}\psi;\\ \mathfrak{M},w\models^{\sigma}\exists x\,\varphi&\text{iff}&\parbox[t][0.0pt][t]{119.50157pt}{$\mathfrak{M},w\models^{\sigma^{\prime}}\varphi$, for some $x$-variant $\sigma^{\prime}$ of $\sigma$;}\\ \mathfrak{M},w\models^{\sigma}\Diamond_{i}\varphi&\text{iff}&\parbox[t][0.0pt][t]{119.50157pt}{$\mathfrak{M},v\models^{\sigma}\varphi$, for some $v\in W$ such that $wR_{i}v$;}\\ \mathfrak{M},w\models^{\sigma}\langle\lambda x.\varphi\rangle(\tau)&\text{iff}&\parbox[t][0.0pt][t]{119.50157pt}{$\tau$ designates under $\sigma$ at $w$ of $\mathfrak{M}$ and $\mathfrak{M},w\models^{\sigma^{\prime}}\varphi$, where $\sigma^{\prime}$ is the $x$-variant of $\sigma$ with $\sigma^{\prime}(x)={\tau}^{\mathfrak{M},{w}}_{\sigma}$.}\end{array}}$
@@ Appendix B Proofs for Section 3
Clearly, for any interpretation $\mathfrak{M}^{\ast}$ with $\mathfrak{M}^{\ast}\models\mathcal{O}^{\ast}$, we have $\mathfrak{M}^{\ast}\models\mathcal{O}$ as well.
@@ Appendix C Proofs for Section 4
$\top$ $\sqsubseteq\exists^{=1}u.\{a\}^{\sharp},\text{ for every }a\in{\sf N_{I}}\text{ in }\mathcal{O},$
@@ D.1 Proof of Theorem 11
$\boldsymbol{w}=w_{0}i_{0}w_{1}\cdots i_{m-1}w_{m},$ (7)
where $1\leq i_{j}\leq n$, $i_{j}\not=i_{j+1}$, and each $R_{i}$ is the smallest equivalence relation containing all pairs of the form $(\boldsymbol{w},\boldsymbol{w}iw)\in W\times W$. We call $w_{0}$ the root of $\mathfrak{F}$. We define the depth of $\mathfrak{F}$ as the maximal $m$ such that $W$ contains a word of the form (7). A tree-shaped $\mathbf{S5}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}$-model takes the form $\mathfrak{M}=(\mathfrak{F},\Delta,\mathcal{I})$ with $\mathfrak{F}$ a tree-shaped $\mathbf{S5}^{n}$-frame.
@@
where $W^{\ast}$ is the set of words $\boldsymbol{w}$ of the form (7) with $(w_{j},w_{j+1})\in R_{i_{j}}$, $w_{j}\not=w_{j+1}$, $i_{j}\not=i_{j+1}$, and $m\leq K$, and $R^{\ast}_{i}$ is the smallest equivalence relation containing all pairs $(\boldsymbol{w},\boldsymbol{w}iw)\in W^{\ast}\times W^{\ast}$. Define a model $\mathfrak{M}^{\ast}=(\mathfrak{F}^{\ast},\Delta^{\ast},\mathcal{I}^{\ast})$ by setting $\Delta^{\ast}=\Delta$ and $\mathcal{I}^{\ast}_{\boldsymbol{w}}=\mathcal{I}_{w_{m}}$ for $\boldsymbol{w}\in W^{\ast}$ of the form (7). Clearly $\mathfrak{M}^{\ast}$ is tree-shaped. Moreover, for all $\smash{\smash{\mathcal{ML}}^{n}_{\smash{\mathcal{ALCO}_{\!u}}}}$-concepts $D$ of modal depth not exceeding $K$, one can show by induction that $D^{\mathcal{I}^{\ast}_{w_{0}}}=D^{\mathcal{I}_{w_{0}}}$. In particular, $C^{\mathcal{I}^{\ast}_{w_{0}}}\neq\emptyset$. ∎
@@
    for every $w\in W$ and every ${\boldsymbol{t}}\in\boldsymbol{q}(w),\text{ with }\{a\}\in{\boldsymbol{t}}$, there exists exactly one $\rho\in\mathfrak{R}$ such that $\{a\}\in\rho(w)$.
@@
For every $C\in\mathsf{con}(C_{0})$, $w\in W$ and $\rho\in\Delta$, we have $\rho\in C^{\mathcal{I}_{w}}$ iff $C\in\rho(w)$.
@@
For each $\boldsymbol{w}\in W$ and ${\boldsymbol{t}}\in\boldsymbol{q}(\boldsymbol{w})$ we fix a proto-run, $\rho_{\boldsymbol{w},{\boldsymbol{t}}}\in\mathfrak{R}$, for $\boldsymbol{w}$ and ${\boldsymbol{t}}$ with $\rho_{\boldsymbol{w},{\boldsymbol{t}}}(\boldsymbol{w})={\boldsymbol{t}}$. Let $\mathcal{F}(\boldsymbol{w})$ denote the set of selected proto-runs through $\boldsymbol{q}(\boldsymbol{w})$, that is $\mathcal{F}(\boldsymbol{w})=\{\rho_{\boldsymbol{w},{\boldsymbol{t}}}\mid{\boldsymbol{t}}\in\boldsymbol{q}(\boldsymbol{w})\}$. We also fix for any $\Diamond_{i}C\in{\boldsymbol{t}}\in\boldsymbol{q}(\boldsymbol{w})$ a proto-witness $\boldsymbol{w}^{\prime}=v_{\boldsymbol{w},i,C,{\boldsymbol{t}}}\in W$ such that $\boldsymbol{w}^{\prime}$ takes the form $\boldsymbol{w}iv$ (and so $(\boldsymbol{w},\boldsymbol{w}^{\prime})\in R_{i}$) and $C\in\rho_{\boldsymbol{w},{\boldsymbol{t}}}(\boldsymbol{w}^{\prime})$, whenever such a $\boldsymbol{w}^{\prime}$ exists. Note that such a $\boldsymbol{w}^{\prime}$ might not exist since by the definition of $R_{i}$ as equivalence closure in tree-shaped models the witness for $\Diamond_{i}C$ might also be ${\boldsymbol{t}}$ itself if $C\in{\boldsymbol{t}}$ or $\boldsymbol{w}$ could take the form $\boldsymbol{w}_{0}iw$ for some $\boldsymbol{w}_{0}$ and either $C\in\rho_{\boldsymbol{w},{\boldsymbol{t}}}(\boldsymbol{w}_{0})$ or there is $w^{\prime}$ with $\boldsymbol{w}_{0}iw^{\prime}\in W$ and $C\in\rho_{\boldsymbol{w},{\boldsymbol{t}}}(\boldsymbol{w}_{0}iw^{\prime})$.
@@
$W_{j+1}^{i}=\{v_{\boldsymbol{w},i,C,{\boldsymbol{t}}}\mid\Diamond_{i}C\in{\boldsymbol{t}}\in\boldsymbol{q}(\boldsymbol{w}),\boldsymbol{w}\in W_{j}\}.$
@@
$R_{i}^{\prime}=R_{i}\cap(W^{\prime}\times W^{\prime}),\quad\boldsymbol{q}^{\prime}(\boldsymbol{w})=\boldsymbol{q}(\boldsymbol{w})\text{ for $\boldsymbol{w}\in W^{\prime}$}$
Also let $\mathfrak{R}^{\prime}$ denote the set of restrictions of the proto-runs $\rho_{\boldsymbol{w},{\boldsymbol{t}}}$ with $\boldsymbol{w}\in W^{\prime}$ to $W^{\prime}$. Let $\mathfrak{F}^{\prime}=(W^{\prime},R_{1}^{\prime},\ldots,R_{n}^{\prime})$ and $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$. Then $\mathfrak{Q}^{\prime}$ is a quasimodel except that the proto-runs in $\mathfrak{R}^{\prime}$ might not all satisfy (R2). Note that (R2) holds for every proto-run $\rho_{\boldsymbol{w},{\boldsymbol{t}}}$ at the world $\boldsymbol{w}$ for those $\Diamond_{i}C$ for which $i_{m-1}\neq i$ (assuming that $\boldsymbol{w}$ takes the form (7)) but that there might not be witnesses for $\Diamond_{i}C$ on $\boldsymbol{w}^{\prime}\neq\boldsymbol{w}$ and if $i_{m-1}=i$. To ensure (R2) also in these cases we introduce, in a careful way, copies of existing worlds that provide the witnesses for such $\Diamond_{i}C$.
@@
Take for any $\boldsymbol{w}\in W^{\prime}$ and $\rho\in\mathfrak{R}^{\prime}$ the bijection $\sigma_{\boldsymbol{w},\rho}\in B(\mathfrak{R}^{\prime})$ that swaps $\rho$ with the proto-run $\rho_{\boldsymbol{w},{\boldsymbol{t}}}$ with ${\boldsymbol{t}}=\rho(\boldsymbol{w})$ and maps the remaining runs to themselves. The $\boldsymbol{w}$-repair set, $\textrm{Rep}(\boldsymbol{w})$, is the set of all $\sigma_{\boldsymbol{w},\rho}$ with $\rho\in\mathfrak{R}^{\prime}$. Note that the identity function, $id$, on $\mathfrak{R}^{\prime}$ is an element of $\textrm{Rep}(\boldsymbol{w})$.
For any $(\boldsymbol{w},\sigma)\in W^{\prime}\times B(\mathfrak{R}^{\prime})$, define
$Suc_{i}(\boldsymbol{w},\sigma)=\{(\boldsymbol{w}^{\prime},\sigma^{\prime})\in W^{\prime}\times B\mid\boldsymbol{w}^{\prime}=\boldsymbol{w}iw\in W^{\prime},\\ \sigma^{\prime}=\tau\circ\sigma\text{ for some }\tau\in\textrm{Rep}(\boldsymbol{w})\}$
@@
$\boldsymbol{u}=(\boldsymbol{w}_{0},\sigma_{0})i_{0}(\boldsymbol{w}_{1},\sigma_{1})i_{1}\cdots i_{m-1}(\boldsymbol{w}_{m},\sigma_{m}),$ (8)
where $(\boldsymbol{w}_{0},\sigma_{0})=(w_{0},id)$ and $(\boldsymbol{w}_{i+1},\sigma_{i+1})\in Suc_{i}(\boldsymbol{w}_{i},\sigma_{i})$ for all $0\leq i<m$.
Define $\boldsymbol{q}^{\prime\prime}$ by setting $\boldsymbol{q}^{\prime\prime}(\boldsymbol{u})=\boldsymbol{q}^{\prime}(\boldsymbol{w}_{m})$ for any $\boldsymbol{u}\in W^{\prime\prime}$ of the form (8). Associate with any run $\rho\in\mathfrak{R}^{\prime}$ a unique run $\rho^{\uparrow}$ on $W^{\prime\prime}$ by setting $\rho^{\uparrow}(\boldsymbol{u}):=\sigma_{m}(\rho)(\boldsymbol{w}_{m})$ for any $\boldsymbol{u}\in W^{\prime\prime}$ of the form (8) and let ′′=$\mathfrak{R}^{\prime\prime}=\{\rho^{\uparrow}\mid\rho\in\mathfrak{R}^{\prime}\}$.
@@
Observe first that $\rho\mapsto\rho^{\uparrow}$ is a bijection from $\mathfrak{R}^{\prime}$ onto $\mathfrak{R}^{\prime\prime}$. To show this observe that the mapping is surjective by definition. To show that it is injective, assume that $\rho_{1}\not=\rho_{2}$. Take $\boldsymbol{w}_{m}\in W^{\prime}$ of the form (7) with $\rho_{1}(\boldsymbol{w}_{m})\neq\rho_{2}(\boldsymbol{w}_{m})$. Then let
$\boldsymbol{v}=(\boldsymbol{w}_{0},id)i_{0}(\boldsymbol{w}_{1},id)i_{1}\cdots i_{m-1}(\boldsymbol{w}_{m},id)\in W^{\prime\prime},$
where $\boldsymbol{w}_{i+1}=\boldsymbol{w}_{i}iw_{i+1}$ for $i<m$. We have $\rho_{1}^{\uparrow}(\boldsymbol{v})=id(\rho_{1})(\boldsymbol{w}_{m})\neq id(\rho_{2})(\boldsymbol{w}_{m})$, as required.
@@
We prove (M1) and (M2) in one go by showing that, for any $\boldsymbol{u}\in W^{\prime\prime}$ of the form (8) and ${\boldsymbol{t}}\in\boldsymbol{q}^{\prime\prime}(\boldsymbol{u})$, there is a bijection between $V_{0}=\{\rho\in\mathfrak{R}^{\prime\prime}\mid\rho(\boldsymbol{u})={\boldsymbol{t}}\}$ and $V_{1}=\{\rho\in\mathfrak{R}^{\prime}\mid\rho(\boldsymbol{w}_{m})={\boldsymbol{t}}\}$ (recall that $\boldsymbol{q}^{\prime\prime}(\boldsymbol{u})=\boldsymbol{q}^{\prime}(\boldsymbol{w}_{m})$).
Indeed, we claim that $F\colon V_{0}\rightarrow V_{1}$ defined by setting $F(\rho^{\uparrow})=\sigma_{m}(\rho)$ is a bijection. As $\rho\mapsto\rho^{\uparrow}$ is bijective and $\rho^{\uparrow}(\boldsymbol{u})=\sigma_{m}(\rho)(\boldsymbol{w}_{m})$, $F$ is a well defined function into $V_{1}$. It is injective since $\sigma_{m}$ is also bijective, by definition. It is surjective since for any $\rho\in\mathfrak{R}^{\prime}$ with $\rho(\boldsymbol{w}_{m})={\boldsymbol{t}}$ we have $F((\sigma_{m}^{-1}(\rho))^{\uparrow})=\sigma_{m}\sigma_{m}^{-}(\rho)=\rho$ and $\sigma_{m}^{-1}(\rho)\in\mathfrak{R}^{\prime}$.
We next show (R1) and (R2). Assume that $\Diamond_{i}C\in\mathsf{con}(C_{0})$ and $\rho^{\uparrow}\in\mathfrak{R}^{\prime\prime}$. Consider $\boldsymbol{u}\in W^{\prime\prime}$ of the form (8).
For (R1), assume there exists $\boldsymbol{v}\in W^{\prime\prime}$ such that $\boldsymbol{u}R_{i}^{\prime\prime}\boldsymbol{v}$ and $C\in\rho^{\uparrow}(\boldsymbol{v})$. We have to show that $\Diamond_{i}C\in\rho^{\uparrow}(\boldsymbol{u})$. Observe that the runs in $\mathfrak{R}^{\prime}$ satisfy (R1) in $\mathfrak{Q}^{\prime}$ and that, moreover,
@@
    if $\Diamond_{i}C\in\rho(\boldsymbol{w})$ with $\rho\in\mathfrak{R}^{\prime}$ and $(\boldsymbol{w},\boldsymbol{w}^{\prime})\in R_{i}^{\prime}$, then $\Diamond_{i}C\in\rho(\boldsymbol{w}^{\prime})$.
@@
    $\boldsymbol{v}=\boldsymbol{u}$. Then $\Diamond_{i}C\in\rho^{\uparrow}(\boldsymbol{u})$ since we have $\Diamond_{i}C\in{\boldsymbol{t}}$ whenever $C\in{\boldsymbol{t}}$ for $\Diamond_{i}C\in\mathsf{con}(C_{0})$ for any ${\boldsymbol{t}}$.
@@
    $\boldsymbol{v}=\boldsymbol{u}i(\boldsymbol{u}iw_{m+1},\sigma_{m+1})$ for some $(w_{m+1},\sigma_{m+1})$ with $\boldsymbol{w}_{m}iw_{m+1}\in W^{\prime}$ with $\sigma_{m+1}=\tau\circ\sigma_{n}$ and $\tau\in\textrm{Rep}(\boldsymbol{w}_{m})$.
    We have $\rho^{\uparrow}(\boldsymbol{u})=\sigma_{m}(\rho)(\boldsymbol{w}_{m})$ and $\rho^{\uparrow}(\boldsymbol{v})=\tau\circ\sigma_{m}(\rho)(\boldsymbol{w}_{m}iw_{m+1})$. From $C\in\tau\circ\sigma_{m}(\rho)(\boldsymbol{w}_{m}iw_{m+1})$, it follows that $\Diamond_{i}C\in\tau\circ\sigma_{m}(\rho)(\boldsymbol{w}_{m})$ since $\tau\circ\sigma_{m}(\rho)$ satisfies condition (R1) for runs in $\mathfrak{Q}^{\prime}$. Now, $\tau\circ\sigma_{m}(\rho)$ coincides with $\sigma_{m}(\rho)$ on $\boldsymbol{w}_{m}$. Hence, $\Diamond_{i}C\in\sigma_{m}(\rho)(\boldsymbol{w}_{m})$, as required.